If n is an integer and 4 is the remainder when 4n - 1 is divided by 5, then n could be
Answer by RAY100(1637)

Let x = Quotient of above
.
Then 5x+4 = 4n-1,,,,, that is quotient times divisor is 4 less than dividend,,,hence remainder 4
.
5x +5 =4n
.
(5x+5) / 4 =n
.
itterate x until n = integer
.
@ x=7,,,n=10
.
check
.
( 4*10 -1) / 5 = 39/5 = 7 rem4